# Dexterous Manipulation Policies from RGB Human Videos via 4D Hand-Object Trajectory Reconstruction

### v1

@@ Abstract
Multi-finger robotic hand manipulation and grasping are challenging due to the high-dimensional action space and the difficulty of acquiring large-scale training data. Existing approaches largely rely on human teleoperation with wearable devices or specialized sensing equipment to capture hand-object interactions, which limits scalability. In this work, we propose VideoManip, a device-free framework that learns dexterous manipulation directly from RGB human videos. Leveraging recent advances in computer vision, VideoManip reconstructs explicit 4D robot-object trajectories from monocular videos by estimating human hand poses, object meshes, and retargets the reconstructed human motions to robotic hands for manipulation learning. To make the reconstructed robot data suitable for dexterous manipulation training, we introduce hand-object contact optimization with interaction-centric grasp modeling, as well as a demonstration synthesis strategy that generates diverse training trajectories from a single video, enabling generalizable policy learning without additional robot demonstrations. In simulation, the learned grasping model achieves a 70.25% success rate across 20 diverse objects using the Inspire Hand. In the real world, manipulation policies trained from RGB videos achieve an average 62.86% success rate across seven tasks using the LEAP Hand, outperforming retargeting-based methods by 15.87%. Project videos are available at videomanip.github.io.
@@ I Introduction
To what extent can dexterous robot manipulation be learned using human video–only supervision? Existing works have trained manipulation policies through human data obtained via hand-held grippers [1], wearables such as smart glasses [2] and headsets [3, 4], or multi-camera studio setups [5]. Although effective, these methods typically require specialized hardware, controlled capture environments, or direct human involvement, which limits scalability. In contrast, RGB videos are ubiquitous and available at scale, offering the potential for scalable learning without specialized hardware. However, extracting accurate supervision for robot–object interactions from such data remains challenging, as these videos lack robot actions and precise 3D information. In this work, we study dexterous manipulation learning from RGB-only human videos and propose a framework that leverages video-only supervision, without requiring robot data or additional sensors. We evaluate our approach under progressively relaxed data conditions, from in-scene human videos to in-the-wild videos.
@@
Prior work has explored leveraging egocentric human videos for manipulation learning; however, these approaches either have not been successfully deployed on multi-fingered robotic hands [6, 7], rely on robot demonstrations to fine-tune models pre-trained on in-the-wild dataset [8, 9, 10], are trained on videos collected in deployment-specific scenarios [11], or assume access to pre-scanned object models [12]. Several studies reconstruct human hand poses or object meshes from images or videos of humans manipulating objects. However, these methods either have not demonstrated effective use of the reconstructed information for learning dexterous robotic hand manipulation [13, 14, 15], or fail to exploit the reconstructed trajectories for direct imitation learning, instead relying on designed reward functions for manipulation policy learning [16] or requiring additional policy fine-tuning [9]. In parallel, another line of research focuses on human teleoperation of robotic hands [17, 18], often requiring head-mounted vision devices or specialized sensing equipment to track human hands [4, 3, 19]. These dependencies on specialized hardware limit scalability and restrict data diversity.
@@
To address these challenges, we propose VideoManip, a framework that learns dexterous grasping and manipulation entirely from RGB human videos, without requiring wearables or external sensor devices, robot demonstrations for post-training, or pre-scanned object models (see Figure LABEL:fig:teaser). Leveraging recent advances in computer vision, such as metric depth estimation and image-to-mesh reconstruction, our system recovers explicit 4D hand–object trajectories from monocular videos by estimating human hand poses, object meshes, and object scales, and subsequently retargets the reconstructed human motions to robotic hand embodiments. For in-scene videos, reconstructed robot trajectories are transformed to the robot frame using known camera–robot calibration. For in-the-wild videos, we estimate the gravity direction from visual observations and use it to align camera-centric trajectories with a physically meaningful world frame.
@@
To make the reconstructed robot data applicable for dexterous manipulation training, we introduce two components that improve the robustness of grasp model training and the generalization of manipulation policies. First, we perform (i) differential hand pose optimization by predicting hand–object contact maps to encourage physically plausible interactions, and (ii) interaction-centric grasp modeling to exploit the optimized contact map for valid grasp learning. Second, we adopt DemoGen [20] to synthesize diverse demonstrations from a single reconstructed video trajectory, enabling one-to-many trajectory generation for generalizable manipulation policy training. We validate our approach on two task categories: object grasping in simulation, and object manipulation in real-world. Comprehensive evaluations and ablations show that the learned grasping model achieves a 70.25% success rate across 20 objects using the Inspire Hand, while the learned manipulation policies with the LEAP Hand achieve an average 62.86% success rate across seven real-world manipulation tasks, including three tasks learned from in-scene human videos and four tasks learned from in-the-wild human videos.
@@
Contributions. (i) VideoManip: a framework for learning multi-fingered robotic hand grasping and manipulation directly from RGB human videos, without requiring additional robot demonstrations, wearable devices, or additional sensing hardware. (ii) An explicit, trainable 4D robot hand-object trajectory reconstruction framework from both in-scene and in-the-wild human videos. (iii) Using reconstructed robot data, we enable dexterous robotic hands to perform diverse object grasping and generalizable manipulation tasks from a single RGB human video per object or task.
@@ II-A Manipulation Learning from Human Videos
More recent systems demonstrate dexterous manipulation without robot demonstrations, but rely on specialized sensing hardware. AINA [2] uses smart glasses to capture hand–object point clouds, while related methods depend on wearable headsets [3, 19] or hand motion capture systems [12, 30], introducing additional hardware dependencies. In contrast, our method operates directly on standard RGB human videos. In parallel, world models [31, 32] and video generative models [27, 33, 34] enable zero-shot manipulation by predicting future object trajectories, primarily for gripper-based tasks. LVP [35] extends this approach to dexterous hands by retargeting predicted human hand motions, but it does not explicitly model contact-rich hand-object interactions and can produce physically infeasible motions. In contrast, our work emphasizes contact-aware grasp modeling and reconstructs reliable, learnable 4D hand-object trajectories from human videos, while remaining complementary to video generation based methods.
@@ II-B Human-Object Interaction Reconstruction from Videos
Reconstructing hands and objects from images and videos is a long-studied research area. Modern approaches leverage powerful models, such as transformers, to infer low-dimensional hand pose and shape representations [36, 37]. With strong 3D supervision, it is feasible to learn shared models across multiple object categories using diverse object representations, including meshes [38], point clouds [39] and primitives [40]. With these advancements, joint reasoning about hand–object interactions from images and videos has become increasingly feasible [41, 13, 42], leveraging methods like diffusion models [43], implicit signed distance fields [44, 45] and differentiable pose estimation to optimize the 3D hand–object contact [46, 47]. Recent works have utilized reconstructed hand–object interactions from images and videos to train dexterous robot grasping and manipulation policies [14, 9, 48], often by explicitly retargeting human hand motions to robot actions [49, 50]. Existing methods, however, either require extensive filtering due to reconstruction inaccuracies [14], incur long reconstruction times [45], or rely on additional robot demonstrations because reconstructed data alone is insufficient [9]. In contrast, VideoManip learns manipulation policies directly from robot trajectories reconstructed from RGB human videos.
@@ III Method
Figure 1: Overview of the VideoManip framework.We first reconstruct 4D robot–object interaction trajectories from RGB human videos via recent advances in 3D vision (Sec. III-A). To utilize the reconstructed data for dexterous grasping and manipulation learning, we perform grasp contact optimization and interaction-centric grasp modeling, and synthesize trajectories for generalizable manipulation (Sec. III-B). Finally, we deploy the trained models for real-world dexterous grasping and manipulation.
@@
Overview. In this section, we present (i) the overall pipeline for reconstructing hand mesh $\mathbf{H}$ and object mesh $\mathbf{O}$ from RGB human videos in Sec III-A, which yields explicit robot hand trajectories, finger-level manipulations, and the resulting object pose changes; and (ii) the learning of grasping and manipulation policies from these reconstructed hand actions and object point clouds in Sec III-B.
@@
Assumptions. (i) The raw dataset consists exclusively of egocentric human videos captured with a static camera, including both in-scene and in-the-wild recordings. Only one video is available for each task or object. (ii) No pre-scanned object meshes, object size information, depth measurements, or camera intrinsics are assumed to be available. For in-scene videos, hand–eye calibration is available, enabling computation of the camera extrinsic matrix to transform reconstructed trajectories from the camera frame to the robot base frame. In contrast, in-the-wild videos are recorded outside robot setups without extrinsic camera calibration; however, the camera is assumed to remain static in both cases. (iii) No robot data or wearable and external sensing devices are used.
III-A 4D Hand-Object Trajectory Reconstruction from Video
@@
Given an RGB video of a human interacting with objects, $\mathcal{V}\in\mathbb{R}^{T\times H\times W\times 3}$, with frames $I_{t}=\mathcal{V}(t)$ for $t=1,\ldots,T$, we leverage off-the-shelf 3D vision techniques to reconstruct the object mesh $\mathcal{O}$, estimate its 6D pose, and recover the human hand pose, followed by robot hand retargeting. To do so, we first adopt MoGe-2 [51] to estimate metric depth maps and camera intrinsics, which define a joint, metric 3D coordinate frame that enables physically consistent hand-object spatial alignment in all subsequent reconstructions.
@@
Object Mesh Reconstruction and Pose Estimation. Our next step is to reconstruct object mesh $\mathcal{O}$ and its 6D pose from the video. Unlike prior manipulation work [2], which tracks only surface point clouds (2.5D), we reconstruct a complete 3D object mesh directly from RGB observations to enable accurate grasp contact modeling. We first identify manipulated objects in the video and obtain object masks using a state-of-the-art video segmentation approach, Segment Anything Model 2 (SAM 2) [52]. The masked regions are then cropped to produce object-centric images and fed into MeshyAI for image-to-mesh generation [53]. While the reconstructed object meshes $\mathcal{O}$ capture the objects’ visual appearance and overall shape, they lack accurate metric scale. and may be larger or smaller than the true objects observed in the video. State-of-the-art pose estimation methods, such as FoundationPose [54], assume access to object meshes with correct real-world dimensions; thus, violating this assumption can lead to erroneous pose estimates.
@@
Human Hand Mesh Estimation and Robot Hand Retargeting. We use the off-the-shelf hand mesh recovery model HaMeR [37] to reconstruct the human hand mesh $\mathcal{H}$ across all frames of the video. HaMeR represents the hand mesh using a low-dimensional parameterization $\mathbf{h}=(\theta,\beta)$, where $\theta$ denotes the hand pose and $\beta$ denotes the hand shape. HaMeR uses a weak-perspective camera model, which is inherently depth-ambiguous and sensitive to focal length errors. To align the reconstructed hand mesh $\mathcal{H}$ within a joint coordinate space with the object, we reuse the metric depth maps predicted by Moge-2, which are consistent with those used for object pose estimation, to ensure that $\mathcal{H}$ and $\mathcal{O}$ share the same depth reference. We then compute the corrected hand depth $t^{\prime}_{z}$ by averaging the metric depths at the 2D keypoints predicted by HaMeR. For every frame $I_{t}$, The reconstructed human hand poses ($\theta_{t}$ and $\beta_{t}$) are then retargeted to the robot hand configuration $q_{t}$, including the wrist pose and finger joint angles, via an optimization that minimizes the error between selected robot link keypoints and their corresponding human hand joints [49]. Given the robot URDF file, the robot mesh $\mathcal{R}$ can be generated for any robot hand configuration $q$.
@@
In-the-Wild Videos Calibration. For trajectories reconstructed from in-scene videos, we can directly transform actions and states into the robot base (world) frame using a calibrated rigid transformation $^{\text{world}}T_{\text{cam}}$. However, this is not feasible for in-the-wild videos, where actions are expressed in the camera coordinate frame and are distorted by unknown camera orientation and pose. As illustrated in Figure 1, before gravity alignment the reconstructed objects (e.g., the bowl and bottle) do not lie on the horizontal plane due to a tilted camera. Policies trained on such data would therefore learn camera-frame–specific actions that do not transfer to the physical world frame. To address this issue, we apply GeoCalib [55], an off-the-shelf single-image camera calibration method that leverages universal visual cues of 3D geometry to estimate camera orientation. From the first video frame, GeoCalib infers the gravity direction in the camera frame and the corresponding rotation $^{\text{grav}}R_{\text{cam}}\in\mathrm{SO}(3)$ that aligns gravity with the negative $z$-axis, $\mathbf{g}_{\text{cam}}=[0,0,-1]^{\top}$. We apply this rotation to all reconstructed meshes $\mathcal{R}$, $\mathcal{O}$, and robot configurations $q$, producing gravity-aligned trajectories and point clouds suitable for policy training. While this does not recover the full camera-to-world transformation $^{\text{world}}T_{\text{cam}}$, the estimated $^{\text{grav}}R_{\text{cam}}$ is sufficient to align trajectories from egocentric, in-the-wild videos to a shared reference plane with in-scene videos (e.g., the robot-mounted table), yielding trajectories that closely match those reconstructed from in-scene recordings.
@@ III-B Dexterous Grasp and Manipulation Learning
To improve the physical validity of hand–object interactions and increase trajectory diversity from a single video, we perform differentiable hand–object contact optimization [47] and adopt DemoGen [20] to synthesize spatially randomized manipulation demonstrations from a single reconstructed human trajectory. Following DemoGen’s skill–motion decomposition, we split the reconstructed trajectory into two stages, grasping and manipulation, each learned by a dedicated policy. The grasp stage spans from the time the hand approaches sufficiently close to the object at time $t_{1}$ to when a stable grasp is achieved at time $t_{2}$, before any object manipulation begins. The manipulation stage covers $t_{2},T]$, where the hand operates on the grasped object. The grasping policy is responsible for learning correct hand–object interactions, whereas the manipulation policy controls the object’s relative translation and rotation along with fine-grained finger motions.
Contact Optimization and Interaction-Centric Grasp Modeling. To address grasp reconstruction failures, during the grasping stage $t_{1},t_{2}]$, we refine the grasps using pre-trained ContactOpt [47]. Based on the current interaction between the reconstructed hand mesh $\mathcal{H}$ and grasp object mesh $\mathcal{O}$, we can compute contact maps $C_{\mathcal{H}}(\mathbf{h})$ and $C_{\mathcal{O}}(\mathbf{h})$, which provide smooth, distance-based contact values for each vertex, measuring the proximity of each vertex to the other mesh. For example, the contact value at an object vertex $v_{\mathcal{O}}^{i}$ can be expressed as
@@
[…]

### v2

# Dexterous Manipulation Policies from RGB Human Videos via 3D Hand-Object Trajectory Reconstruction

@@ Abstract
Multi-finger robotic hand manipulation and grasping are challenging due to the high-dimensional action space and the difficulty of acquiring large-scale training data. Existing approaches largely rely on human teleoperation with wearable devices or specialized sensing equipment to collect robot demonstrations, limiting scalability. In this work, we propose VideoManip, a device-free framework that learns dexterous multi-finger manipulation directly from RGB human videos. Leveraging recent advances in computer vision, VideoManip reconstructs explicit robot-object trajectories from monocular videos by estimating human hand poses, object meshes, and retargets the reconstructed human motions to robot hands for manipulation learning. To make the reconstructed robot data suitable for dexterous manipulation training, we introduce hand-object contact optimization with interaction-centric grasp modeling, as well as a demonstration synthesis strategy that generates diverse training trajectories from a single video, enabling generalizable policy learning without additional robot demonstrations. In simulation, the learned grasping model achieves a 70.25% success rate across 20 diverse objects using the Inspire Hand. In the real world, manipulation policies trained from RGB videos achieve an average 62.86% success rate across seven tasks using the LEAP Hand, outperforming retargeting-based methods by 15.87%. Project videos are available at videomanip.github.io.
@@ I Introduction
To what extent can dexterous robot manipulation be learned using human video-only supervision? Existing works have trained manipulation policies through human data obtained via hand-held grippers [1], wearables such as smart glasses [2] and headsets [3, 4], or multi-camera studio setups [5]. Although effective, these methods typically require specialized hardware, controlled capture environments, or direct human involvement, limiting scalability. In contrast, RGB videos are ubiquitous and available at internet scale, offering the potential for scalable learning without specialized hardware. However, extracting accurate supervision for robot-object interactions from such data remains challenging, as these videos lack robot actions and precise 3D information. In this work, we propose a framework for learning dexterous manipulation from RGB-only human videos alone. We evaluate this framework under progressively relaxed data conditions, from in-scene videos to in-the-wild videos.
@@
Prior work explored leveraging egocentric human videos for manipulation learning; however, these approaches either have not been successfully deployed on multi-fingered robotic hands [6, 7], rely on robot demonstrations to fine-tune models pre-trained on in-the-wild dataset [8, 9, 10], are trained on videos collected in deployment-specific scenarios [11], or assume access to pre-scanned object models [12]. Several studies reconstruct human hand poses or object meshes from images or videos of humans manipulating objects. However, these methods either have not demonstrated effective use of the reconstructed information for learning dexterous robotic hand manipulation [13, 14, 15], or fail to exploit the reconstructed trajectories for direct imitation learning, instead relying on designed reward functions for manipulation policy learning [16] or requiring additional policy fine-tuning [9]. In parallel, another line of research focuses on human teleoperation of robotic hands [17, 18], often requiring head-mounted vision devices or specialized sensing equipment to track human hands [4, 3, 19]. These dependencies on specialized hardware limit scalability and restrict data diversity.
@@
To address these challenges, we propose VideoManip, a framework that learns dexterous grasping and manipulation entirely from RGB human videos, without requiring wearables or external sensor devices, robot demonstrations for post-training, or pre-scanned object meshes (Fig. LABEL:fig:teaser). Leveraging recent advances in computer vision, such as metric depth estimation and image-to-mesh reconstruction, VideoManip recovers explicit 3D hand-object trajectories from monocular videos by estimating human hand poses, object meshes, and object scales, and subsequently retargets the reconstructed human motions to robot hands. For in-scene videos, we assume known camera-robot extrinsic calibration and use it to transform reconstructed trajectories from the camera frame to the robot base frame for policy training. In contrast, for in-the-wild videos, which are recorded outside robot setups without extrinsic calibration, we estimate the gravity direction from visual observations and align camera-centric trajectories with a physically meaningful world frame.
@@
To make the reconstructed robot data applicable for dexterous manipulation training, we introduce two components that improve the of grasp-model training and the generalization of manipulation policies. First, we perform (i) differential hand pose optimization by predicting hand-object contact maps to encourage physically plausible interactions, and (ii) interaction-centric grasp modeling to exploit the optimized contact map for valid grasp learning. Second, we leverage DemoGen [20] to synthesize diverse demonstrations from a single reconstructed video trajectory, enabling one-to-many trajectory generation for generalizable manipulation policy training. We evaluate VideoManip on two task categories: object grasping in simulation, and object manipulation in real-world. Comprehensive evaluations and ablations show that the learned grasp model achieves a 70.25 % success rate across 20 objects using the Inspire Hand, while the learned manipulation policies with the LEAP Hand achieve an average 62.86 % success rate across seven real-world manipulation tasks, including three tasks learned from in-scene videos and four tasks learned from in-the-wild videos.
@@
Contributions. (i) VideoManip: a framework for learning multi-fingered robotic hand grasping and manipulation directly from RGB human videos alone, without requiring robot demonstrations, wearable devices, or external sensing hardware; (ii) An explicit, trainable 3D robot hand-object trajectory reconstruction pipeline from both in-scene and in-the-wild human videos; (iii) Using reconstructed robot data, we enable dexterous robot hands to perform diverse object grasping and generalizable manipulation tasks from a single RGB human video per object or task.
@@ II-A Manipulation Learning from Human Videos
Another line of works demonstrate dexterous manipulation without robot demonstrations, but rely on specialized sensing hardware. AINA [2] uses smart glasses to capture hand–object point clouds, while related methods depend on wearable headsets [3, 19] or hand motion capture systems [12, 30], introducing additional hardware dependencies. In parallel, world models [31, 32] and video generative models [27, 33, 34] enable zero-shot manipulation by predicting future object trajectories, primarily for gripper-based tasks. LVP [35] extends this approach to dexterous hands by retargeting predicted human hand motions, but it does not explicitly model hand-object contacts and can produce infeasible motions. In contrast, our work emphasizes contact-aware grasp modeling and reconstructs reliable, learnable 3D hand-object trajectories from human videos, while remaining complementary to video generation based methods.
@@ II-B Human-Object Interaction Reconstruction from Videos
Reconstructing hands and objects from images and videos is a long-studied research area. Modern approaches leverage powerful models, such as transformers, to infer low-dimensional hand pose and shape representations [36, 37]. With strong 3D supervision, it is feasible to learn shared models across multiple object categories using diverse object representations, including meshes [38], point clouds [39] and primitives [40]. With these advancements, joint reasoning about hand–object interactions from images and videos has become increasingly feasible [41, 13, 42], leveraging methods like diffusion models [43], implicit signed distance fields [44, 45] and differentiable pose estimation to optimize the 3D hand-object contact [46, 47]. Recent works have utilized reconstructed hand-object interactions from images and videos to train dexterous robot grasping and manipulation policies [14, 9, 48], often by explicitly retargeting human hand motions to robot actions [49, 50]. Existing methods, however, either require extensive filtering due to reconstruction inaccuracies [14], incur long reconstruction times [45], or rely on additional robot demonstrations because reconstructed data alone is insufficient [9]. In contrast, VideoManip learns manipulation policies directly from robot trajectories reconstructed from RGB human videos.
@@ III Method
Figure 1: Overview of the VideoManip framework. We first reconstruct 3D robot–object interaction trajectories from RGB human videos via recent advances in 3D vision (Sec. III-A). To utilize the reconstructed data for dexterous grasping and manipulation learning, we perform grasp contact optimization and interaction-centric grasp modeling, and synthesize trajectories for generalizable manipulation (Sec. III-B). Finally, we deploy the trained models for real-world dexterous grasping and manipulation.
@@
In this section, we present (i) the overall pipeline for reconstructing hand mesh $\mathbf{H}$ and object mesh $\mathbf{O}$ from RGB human videos in Sec III-A, which yields explicit 3D robot hand trajectories, finger-level manipulations, and the resulting object pose changes; and (ii) the learning of grasping and manipulation policies from these reconstructed hand actions and object point clouds in Sec III-B.
@@
Assumptions. (i) The raw dataset consists exclusively of egocentric human videos captured with a static camera, including both in-scene and in-the-wild recordings. Only one video is available for each task or object. (ii) No pre-scanned object meshes, object size information, depth measurements, or camera intrinsics are assumed to be available. (iii) No robot data or wearable and external sensing devices are used.
III-A 3D Hand-Object Trajectory Reconstruction from Video
@@
Given an RGB video of a human interacting with objects, $\mathcal{V}=\{I_{1},\ldots,I_{T}\}$, with image frames $I_{t}\in\mathbb{R}^{H\times W\times 3}$, we leverage existing 3D vision techniques to reconstruct the object mesh $\mathcal{O}$, estimate its 6D pose, and recover the human hand pose, followed by robot hand retargeting. To do so, we first adopt MoGe-2 [51] to estimate metric depth maps and camera intrinsics, which define a joint, metric 3D coordinate frame that enables physically consistent hand-object spatial alignment in all subsequent reconstructions.
@@
Object Mesh Reconstruction and Pose Estimation. Our next step is to reconstruct object mesh $\mathcal{O}$ and its 6D pose from the video. We reconstruct a complete 3D object mesh directly from RGB observations to enable accurate grasp contact modeling. We first identify manipulated objects in the video and obtain object masks using Segment Anything Model 2 (SAM 2) [52]. We crop the masked regions to produce object-centric images and feed them into MeshyAI for image-to-mesh generation [53]. While the reconstructed object meshes $\mathcal{O}$ capture the objects’ visual appearance and overall shape, they lack accurate metric scale.Pose estimation methods such as FoundationPose [54] assume access to object meshes with accurate real-world dimensions; violating this assumption leads to erroneous pose estimates.
@@
Human Hand Mesh Estimation and Robot Hand Retargeting. To extract robot actions from video, we first recover human hand motion and then retarget it to the robot hand. We use the hand mesh recovery model HaMeR [37] to reconstruct the hand mesh $\mathcal{H}$ for all video frames. HaMeR represents the hand mesh using a low-dimensional parameterization $\mathbf{h}=(\theta,\beta)$, where $\theta$ denotes the hand pose and $\beta$ denotes the hand shape. HaMeR uses a weak-perspective camera model, which is inherently depth-ambiguous and sensitive to focal length errors. To align the reconstructed hand mesh $\mathcal{H}$ within a joint coordinate space with the object, we reuse the metric depth maps predicted by MoGe-2, which are consistent with those used for object pose estimation, to ensure that $\mathcal{H}$ and $\mathcal{O}$ share the same depth reference. We then compute the corrected hand depth $t^{\prime}_{z}$ by averaging the metric depths at the 2D keypoints predicted by HaMeR. For every frame $I_{t}$, The reconstructed human hand poses ($\theta_{t}$ and $\beta_{t}$) are then retargeted to the robot hand configuration $q_{t}$, including the wrist pose and finger joint angles, via an optimization that minimizes the error between selected robot link keypoints and their corresponding human hand joints [49]. Given the robot URDF file, the robot mesh $\mathcal{R}$ can be generated for any robot hand configuration $q$.
@@
In-the-Wild Videos Calibration. Trajectories from in-the-wild videos, unlike those from in-scene videos that can be transformed to the robot base using a calibrated transform $^{\text{world}}T_{\text{cam}}$, lack this transformation. Their actions are expressed in the camera frame and are misaligned with the world frame due to unknown camera orientation and pose. As illustrated in Figure 1, before gravity alignment the reconstructed objects (e.g., the bowl and bottle) do not lie on the horizontal plane due to a tilted camera. Policies trained on such data would therefore learn camera-frame specific actions that do not transfer to the physical world frame. To address this issue, we apply GeoCalib [55], a single-image camera calibration method that leverages universal visual cues of 3D geometry to estimate camera orientation. From the first video frame, GeoCalib infers the gravity direction in the camera frame and the corresponding rotation $^{\text{grav}}R_{\text{cam}}\in\mathrm{SO}(3)$ that aligns gravity with the negative $z$-axis, $\mathbf{g}_{\text{cam}}=[0,0,-1]^{\top}$. We apply this rotation to all reconstructed meshes $\mathcal{R}$, $\mathcal{O}$, and robot configurations $q$, producing gravity-aligned trajectories and point clouds suitable for policy training. While this does not recover the full camera-to-world transformation $^{\text{world}}T_{\text{cam}}$, the estimated $^{\text{grav}}R_{\text{cam}}$ is sufficient to align trajectories from egocentric, in-the-wild videos to a shared reference plane with in-scene videos (e.g., the robot-mounted table), yielding trajectories that closely match those reconstructed from in-scene recordings.
@@ III-B Dexterous Grasp and Manipulation Learning
To improve the physical validity of hand-object interactions and increase trajectory diversity from a single video, we perform differentiable hand–object contact optimization [47] and adopt DemoGen [20] to synthesize spatially randomized manipulation demonstrations from a single reconstructed human trajectory. Following DemoGen’s skill–motion decomposition, we split the reconstructed trajectory into two stages, grasping and manipulation, each learned by a dedicated policy. The grasp stage spans from the time the hand approaches sufficiently close to the object at time $t_{1}$ to when a stable grasp is achieved at time $t_{2}$, before any object manipulation begins. The manipulation stage covers $t_{2},T]$, where the hand operates on the grasped object. The grasping policy is responsible for learning correct hand–object interactions, whereas the manipulation policy controls the object’s relative translation and rotation along with fine-grained finger motions.
Contact Optimization and Interaction-Centric Grasp Modeling. To address grasp reconstruction failures, during the grasping stage $t_{1},t_{2}]$, we refine the grasps using pre-trained ContactOpt [47].The hand parameters $\mathbf{h}$ determine the geometry of the reconstructed hand mesh $\mathcal{H}$ relative to the object mesh $\mathcal{O}$. We therefore compute hand pose-dependent contact maps $C_{\mathcal{H}}(\mathbf{h})$ and $C_{\mathcal{O}}(\mathbf{h})$, which assign smooth, distance-based contact values to each vertex based on its proximity to the other mesh. For example, the contact value at an object vertex $v_{\mathcal{O}}^{i}$ can be expressed as
@@
$E(\mathbf{h})=\lvert C_{\mathcal{O}}(\mathbf{h})-\hat{C}_{\mathcal{O}}\rvert+\lvert C_{\mathcal{H}}(\mathbf{h})-\hat{C}_{\mathcal{H}}\rvert$.